6/x + x/2 = 4
What is x?

Answer by mukhopadhyay(490)

6/x + x/2 = 4
=> (12+x^2)/2x = 4
=> x^2+12 = 8x
=> x^2-8x+12 = 0
=> (x-6)(x-2) = 0
=> (x-6)=0 or (x-2)=0
=> x=6 or x=2